the demand curve faced by a Monopolist is p= 120-3Q.the marginal cost curve in factory 1and factory 2 are respectively as follows MC1=10+20Q1 MC2=60+5Q2 what is the equilibrium price

1
Expert's answer
2022-02-20T16:00:49-0500

Equilibrium Price

"P=120-3Q"

"MC1=10+20Q1"

"MC2=60+5Q2"

To get the firms quantity at equilibrium from the two factories Marginal Cost we get the total Marginal Cost"(MC_T)" . Equating "MC_T=MR_T" gives the quantity of the firm. "(MR_T)" is equal to the firms AV with similar intercept but double the slope.

"MC_T=Q=Q1+Q2"

"=10+\frac{MC1}{20}+60+\frac{MC2}{5}"

"=70+\frac{5}{20}MC_T=\frac{20}{5}Q+70=4Q+70"

"MR_T=AR\ with\ double\ the\ slope"

"AR=P=120-3Q"

"MR_T=120-6Q"

"MC_T=MR_T;"

"4Q+70=120-6Q"

"10Q=50"

"Q=5"

Replacing the quantity amount in the demand function we get the equilibrium price as;

"P=120-3(5)"

"\bold{P=105}"